# Nonparametric Estimation of Joint Entropy through Partitioned Sample-Spacing Method

### v1

@@ 8.1 Proof of lemma 1
Since $F$ is continuous, $X_{(i+m)}-X_{(i)}\xrightarrow{a.s.}0$. Similarly, one can show that $X_{(i-m)}-X_{(i)}\xrightarrow{a.s.}0$. Since $\xi_{i}=\frac{1}{2m}\sum\limits_{m^{\prime}=-m}^{m-1}X_{(i+m^{\prime})}$,
@@ 8.2 Proof of Theorem 1
Now, using the results from Lemma 1 regarding $X_{(i-m)}$ and $X_{(i+m)}$:
Since $X_{(i-m)}\xrightarrow{a.s.}X_{(i)}$ and $X_{(i)}\xrightarrow{a.s.}x$, by transitivity of almost sure convergence, $X_{(i-m)}\xrightarrow{a.s.}x$. Similarly, since $X_{(i+m)}\xrightarrow{a.s.}X_{(i)}$ and $X_{(i)}\xrightarrow{a.s.}x$, then $X_{(i+m)}\xrightarrow{a.s.}x$.
@@ 8.3 Proof of Proposition 1
where $\Delta x_{i}^{k}=x_{(i+m_{k})}^{k}-x_{(i-m_{k})}^{k}$ and $\Delta y_{j}^{k}=y_{(j+m_{k})}^{k}-y_{(j-m_{k})}^{k}$. The total integral is the sum of integrals over each partition. The integral over a single partition $P_{k}$ can be split into contributions from its $(n_{k}-1)^{2}$ interior cells and its boundary cells: $\int_{C_{k}}\hat{f}_{n,\ell}=InteriorMass_{k}+EdgeMass_{k}$. For any interior cell, the ratio of the grid spacing to the m-spacing simplifies. Using the definition $\eta_{j}^{k}=\frac{1}{2m_{k}}\sum_{l=j-m_{k}}^{j+m_{k}-1}y_{(l)}^{k}$, the difference $\eta_{j+1}^{k}-\eta_{j}^{k}$ becomes a telescoping sum:
@@
The edge mass depends on boundary terms $E_{x}^{k}:=\frac{\xi_{1}^{k}-\xi_{0}^{k}}{\Delta x_{0}^{k}}+\frac{\xi_{n_{k}+1}^{k}-\xi_{n_{k}}^{k}}{\Delta x_{n_{k}}^{k}}$ and $E_{y}^{k}:=\frac{\eta_{1}^{k}-\eta_{0}^{k}}{\Delta y_{0}^{k}}+\frac{\eta_{n_{k}+1}^{k}-\eta_{n_{k}}^{k}}{\Delta y_{n_{k}}^{k}}$. Using the definitions of $\xi_{i}^{k}$ and the boundary conditions on order statistics, we can show $0\leq E_{x}^{k},E_{y}^{k}\leq 1$. The first term of $E_{x}^{k}$ is bounded as:
@@ 8.4 Proof of lemma 2
Convergence of the first term: Let $U_{(j)}^{k}=F_{X|C_{k}}(X_{(j)}^{k})$. These are the order statistics from a Uniform(0,1) distribution, based on the $n_{k}$ samples in $C_{k}$. Thus, $P_{n}^{k}=U_{(i+m_{k})}^{k}-U_{(i-m_{k})}^{k}\sim\mathrm{Beta}(2m_{k},n_{k}-2m_{k}+1)$. Let $\mu_{n}^{k}$ be expectation $\mathbb{E}[P_{n}^{k}]=\frac{2m_{k}}{n_{k}+1}$. Then,
@@
for some $X^{\prime}_{k}\in(X_{(i-m_{k})}^{k},X_{(i+m_{k})}^{k})$. We need to show $X^{\prime}_{k}\xrightarrow{a.s.}x$. The interval $(\xi_{i}^{k},\xi_{i+1}^{k}]$, in which $x$ lies, is defined using averages of order statistics $X_{(j)}^{k}$ partition $k$, similar to $\xi_{i}$ in Lemma 1. Applying the logic of Lemma 1 to the $n_{k}$ samples within partition $k$: $X_{(i\pm m_{k})}^{k}\xrightarrow{a.s.}X_{(i)}^{k}$ (since $m_{k}/n_{k}\to 0$). Furthermore, for $x\in(\xi_{i}^{k},\xi_{i+1}^{k}]$, $X_{(i)}^{k}$ is the relevant order statistic that converges almost surely to $x$. Since $X^{\prime}_{k}$ is between $X_{(i-m_{k})}^{k}$ and $X_{(i+m_{k})}^{k}$, both of which converge almost surely to $x$, it follows by the Sandwich Theorem that $X^{\prime}_{k}\xrightarrow{a.s.}x$. Given that $f_{X|C_{k}}$ is continuous at $x$, $f_{X|C_{k}}(X^{\prime}_{k})\xrightarrow{a.s.}f_{X|C_{k}}(x)$ by the Continuous Mapping Theorem. Finally, the product of the two terms converges almost surely:
@@
8.5 Proof of Theorem 2
Let the partition $k$ containing the point $(x,y)$ be defined by the rectangular region $x_{\text{min},k},x_{\text{max},k}]\times[y_{\text{min},k},y_{\text{max},k}]$. Let $\Delta x_{k}=x_{\text{max},k}-x_{\text{min},k}$ be the width and $\Delta y_{k}=y_{\text{max},k}-y_{\text{min},k}$ be the height of this partition, so its area is $A_{k}=\Delta x_{k}\Delta y_{k}$. The estimator for $f(x,y)$ at the point $(x,y)$ is: $\hat{f}_{n,l}(x,y)=\frac{n_{k}}{n}\cdot\hat{f}_{X}^{k}(x)\cdot\hat{f}_{Y}^{k}(y)$. We aim to show that $\hat{f}_{n,l}(x,y)\xrightarrow{a.s.}f(x,y)$ as $n,\ell\to\infty$.
@@ 8.5 Proof of Theorem 2
By the Mean Value Theorem for Integrals, There exists $\tilde{y}_{k}\in[y_{\text{min},k},y_{\text{max},k}]$ such that $\int_{y_{\text{min},k}}^{y_{\text{max},k}}f(x,s)\,ds=f(x,\tilde{y}_{k})\Delta y_{k}$. Also, there exists $\tilde{x}_{k}\in[x_{\text{min},k},x_{\text{max},k}]$ such that $\int_{x_{\text{min},k}}^{x_{\text{max},k}}f(u,y)\,du=f(\tilde{x}_{k},y)\Delta x_{k}$. Lastly, there exists $(\bar{x}_{k},\bar{y}_{k})\in[x_{\text{min},k},x_{\text{max},k}]\times[y_{\text{min},k},y_{\text{max},k}]$ such that $P(C_{k})=\iint_{k}f(u,s)\,du\,ds=f(\bar{x}_{k},\bar{y}_{k})A_{k}=f(\bar{x}_{k},\bar{y}_{k})\Delta x_{k}\Delta y_{k}$. Substituting these into the expression for $Q_{k}(x,y)$:
@@
Therefore, since $f(x,y)>0$:
$\lim_{n\to\infty}Q_{k}(x,y)=\frac{f(x,y)\cdot f(x,y)}{f(x,y)}=f(x,y)$
@@ 8.7 Proof of Theorem 3
and analogously for $S_{y}$ and $T_{y}$. Then, $\log\hat{f}_{n,\ell}=\log(\frac{n_{k}}{n})+\log S_{x}+\log T_{x}+\log S_{y}+\log T_{y}$. By the mean value theorem inside marginal partition interval $I_{y}^{(k)}$ and $I_{x}^{(k)}$ we have, $T_{x}=\frac{|I_{y}^{(k)}|}{p_{k}}\,f(X^{\prime},y_{k}^{*}),T_{y}=\frac{|I_{x}^{(k)}|}{p_{k}}\,f(x_{k}^{*},Y^{\prime})$ for some $X^{\prime},Y^{\prime},y_{k}^{*},x_{k}^{*}\in C_{k}$ where $p_{k}=\mathbb{P}\{X\in C_{k}\}$. Using $\log(n_{k}/(np_{k}))+\log p_{k}=\log(n_{k}/n)$ and $|C_{k}|=|I_{x}^{(k)}||I_{y}^{(k)}|$ we obtain,
@@
By Hölder’s inequality, for any real numbers $x_{1},\dots,x_{m}$ and any $\varepsilon>0$ we have $\left|\sum_{j=1}^{m}x_{j}\right|^{1+\varepsilon}\leq m^{\varepsilon}\sum_{j=1}^{m}|x_{j}|^{1+\varepsilon}$. Thus,
@@
Similarly $\mathbb{E}[|U_{k}|^{2(1+\delta)}\mid C_{k}]\leq C(np_{k})^{-(1+\delta)}$. Using the inequality for $|u|\leq 1/2$, $|\log(1+u)|\leq C(|u|+u^{2})$, we can bound $\mathbb{E}[|\log{\frac{n_{k}}{n}}|^{1+\delta}]$ as below:
@@
As $\ell^{2}=o(n)$, the first term goes to 0. Also the second term goes to 0. Thus, $\sup_{n}\mathbb{E}[|\log(n_{k}/np_{k})|^{1+\delta}]<\infty$.
@@
We prove this for the $x$ dimension; the proof for the $y$ dimension is identical. Let $P_{x}^{k}=F_{X|C_{k}}(x_{(i+m_{k})}^{k})-F_{X|C_{k}}(x_{(i-m_{k})}^{k})$. This random variable follows a Beta distribution, $P_{x}^{k}\sim\mathrm{Beta}(2m_{k},n_{k}-2m_{k}+1)$. The conditional mean and variance of its logarithm are given in Johnson:
@@
Let $g(u):=|\log f(u)|^{1+\delta}$ with $\int gf<\infty$. Define the upper step function $s_{n}(u):=\sum_{k}\big(\sup_{v\in C_{k}}g(v)\big)\mathbf{1}_{C_{k}}(u)$. Since the mesh $\to 0$ and $f$ is continuous with $f>0$ a.e., $s_{n}\downarrow g$ pointwise and $\int s_{n}f\downarrow\int gf<\infty$. Because $(X^{\prime},y_{k}^{*})\in C_{k}$ and
@@
8.8 Proof Sketch of Theorem 4
@@
8.9 Proof Sketch of Theorem 5
@@
$\boxed{\ \log\hat{f}_{n,\ell}(X)=\underbrace{\log\tfrac{n_{k}}{np_{k}}}_{\text{(A)}}+\underbrace{\sum_{j=1}^{d}\log S_{j}}_{\text{(B)}}+\underbrace{\sum_{j=1}^{d}\log f(\tilde{X}_{j})}_{\text{(C)}}+\underbrace{(d-1)\log\tfrac{|C_{k}|}{p_{k}}}_{\text{(D)}}\ }$

### v2

@@ Appendix A Proof of Lemma 1
Since $F$ is continuous, $X_{(i+m)}-X_{(i)}\xrightarrow{a.s.}0$. Similarly, $X_{(i-m)}-X_{(i)}\xrightarrow{a.s.}0$.
Since $\xi_{i}=\frac{1}{2m}\sum\limits_{m^{\prime}=-m}^{m-1}X_{(i+m^{\prime})}$,
@@ Appendix B Proof of Theorem 2
Now, from Lemma 1, $X_{(i-m)}\xrightarrow{a.s.}X_{(i)}$ and $X_{(i)}\xrightarrow{a.s.}x$. By transitivity of almost sure convergence, $X_{(i-m)}\xrightarrow{a.s.}x$. Similarly, since $X_{(i+m)}\xrightarrow{a.s.}X_{(i)}$ and $X_{(i)}\xrightarrow{a.s.}x$, then $X_{(i+m)}\xrightarrow{a.s.}x$.
@@ Appendix C Proof of Proposition 3
where $\Delta x_{i}^{k}=x_{(i+m_{k})}^{k}-x_{(i-m_{k})}^{k}$ and $\Delta y_{j}^{k}=y_{(j+m_{k})}^{k}-y_{(j-m_{k})}^{k}$. The total integral is the sum of integrals over each partition. The integral over a single partition $P_{k}$ can be split into contributions from its $(n_{k}-1)^{2}$ interior cells and its boundary cells:
$\int_{C_{k}}\hat{f}_{n,\ell}=InteriorMass_{k}+EdgeMass_{k}$. For any interior cell, the ratio of the grid spacing to the m-spacing simplifies. Using the definition $\eta_{j}^{k}=\frac{1}{2m_{k}}\sum_{l=j-m_{k}}^{j+m_{k}-1}y_{(l)}^{k}$, the difference $\eta_{j+1}^{k}-\eta_{j}^{k}$ becomes a telescoping sum:
@@
The edge mass depends on boundary terms $E_{x}^{k}:=\frac{\xi_{1}^{k}-\xi_{0}^{k}}{\Delta x_{0}^{k}}+\frac{\xi_{n_{k}+1}^{k}-\xi_{n_{k}}^{k}}{\Delta x_{n_{k}}^{k}}$ and
$E_{y}^{k}:=\frac{\eta_{1}^{k}-\eta_{0}^{k}}{\Delta y_{0}^{k}}+\frac{\eta_{n_{k}+1}^{k}-\eta_{n_{k}}^{k}}{\Delta y_{n_{k}}^{k}}$. Using the definitions of $\xi_{i}^{k}$ and the boundary conditions on order statistics, we can show $0\leq E_{x}^{k},E_{y}^{k}\leq 1$. The first term of $E_{x}^{k}$ is bounded as:
@@ Appendix D Proof of Lemma 4
Convergence of the first term: Let $U_{(j)}^{k}=F_{X|C_{k}}(X_{(j)}^{k})$. These are the order statistics from a Uniform(0,1) distribution, based on the $n_{k}$ samples in $C_{k}$. Thus,
$P_{n}^{k}=U_{(i+m_{k})}^{k}-U_{(i-m_{k})}^{k}\sim\mathrm{Beta}(2m_{k},n_{k}-2m_{k}+1)$. Let $\mu_{n}^{k}$ be expectation $\mathbb{E}[P_{n}^{k}]=\frac{2m_{k}}{n_{k}+1}$. Then,
@@
for some $X^{\prime}_{k}\in(X_{(i-m_{k})}^{k},X_{(i+m_{k})}^{k})$. We need to show $X^{\prime}_{k}\xrightarrow{a.s.}x$. The interval $(\xi_{i}^{k},\xi_{i+1}^{k}]$, in which $x$ lies, is defined using averages of order statistics $X_{(j)}^{k}$ partition $k$, similar to $\xi_{i}$ in Lemma 1. Applying the logic of Lemma 1 to the $n_{k}$ samples within partition $k$:
$X_{(i\pm m_{k})}^{k}\xrightarrow{a.s.}X_{(i)}^{k}$ (since $m_{k}/n_{k}\to 0$). Furthermore, for $x\in(\xi_{i}^{k},\xi_{i+1}^{k}]$, $X_{(i)}^{k}$ is the relevant order statistic that converges almost surely to $x$. Since $X^{\prime}_{k}$ is between $X_{(i-m_{k})}^{k}$ and $X_{(i+m_{k})}^{k}$, both of which converge almost surely to $x$, it follows by the Sandwich Theorem that
$X^{\prime}_{k}\xrightarrow{a.s.}x$. Given that $f_{X|C_{k}}$ is continuous at $x$, $f_{X|C_{k}}(X^{\prime}_{k})\xrightarrow{a.s.}f_{X|C_{k}}(x)$ by the Continuous Mapping Theorem. Finally, the product of the two terms converges almost surely:
@@
Appendix E Proof of Theorem 5
Let the partition $k$ containing the point $(x,y)$ be defined by the rectangular region
$x_{\text{min},k},x_{\text{max},k}]\times[y_{\text{min},k},y_{\text{max},k}]$. Let $\Delta x_{k}=x_{\text{max},k}-x_{\text{min},k}$ be the width and
$\Delta y_{k}=y_{\text{max},k}-y_{\text{min},k}$ be the height of this partition, so its area is $A_{k}=\Delta x_{k}\Delta y_{k}$. The estimator for $f(x,y)$ at the point $(x,y)$ is: $\hat{f}_{n,l}(x,y)=\frac{n_{k}}{n}\cdot\hat{f}_{X}^{k}(x)\cdot\hat{f}_{Y}^{k}(y)$. We aim to show that $\hat{f}_{n,l}(x,y)\xrightarrow{a.s.}f(x,y)$ as $n,\ell\to\infty$.
@@ Appendix E Proof of Theorem 5
By the Mean Value Theorem for Integrals, There exists $\tilde{y}_{k}\in[y_{\text{min},k},y_{\text{max},k}]$ such that
$\int_{y_{\text{min},k}}^{y_{\text{max},k}}f(x,s)\,ds=f(x,\tilde{y}_{k})\Delta y_{k}$. Also, there exists $\tilde{x}_{k}\in[x_{\text{min},k},x_{\text{max},k}]$ such that
$\int_{x_{\text{min},k}}^{x_{\text{max},k}}f(u,y)\,du=f(\tilde{x}_{k},y)\Delta x_{k}$. Lastly, there exists $(\bar{x}_{k},\bar{y}_{k})\in[x_{\text{min},k},x_{\text{max},k}]\times[y_{\text{min},k},y_{\text{max},k}]$ such that $P(C_{k})=\iint_{k}f(u,s)\,du\,ds=f(\bar{x}_{k},\bar{y}_{k})A_{k}=f(\bar{x}_{k},\bar{y}_{k})\Delta x_{k}\Delta y_{k}$. Substituting these into the expression for $Q_{k}(x,y)$:
@@
Therefore, since $f(x,y)>0$: $\lim_{n\to\infty}Q_{k}(x,y)=\frac{f(x,y)\cdot f(x,y)}{f(x,y)}=f(x,y)$
@@ Appendix G Proof of Theorem 8
and analogously for $S_{y}$ and $T_{y}$. Then, $\log\hat{f}_{n,\ell}=\log(\frac{n_{k}}{n})+\log S_{x}+\log T_{x}+\log S_{y}+\log T_{y}$. By the mean value theorem inside marginal partition interval $I_{y}^{(k)}$ and $I_{x}^{(k)}$ we have,
$T_{x}=\frac{|I_{y}^{(k)}|}{p_{k}}\,f(X^{\prime},y_{k}^{*}),T_{y}=\frac{|I_{x}^{(k)}|}{p_{k}}\,f(x_{k}^{*},Y^{\prime})$ for some $X^{\prime},Y^{\prime},y_{k}^{*},x_{k}^{*}\in C_{k}$ where
$p_{k}=\mathbb{P}\{X\in C_{k}\}$. Using $\log(n_{k}/(np_{k}))+\log p_{k}=\log(n_{k}/n)$ and $|C_{k}|=|I_{x}^{(k)}||I_{y}^{(k)}|$ we obtain,
@@
By Hölder’s inequality, for any real numbers $x_{1},\dots,x_{m}$ and any $\varepsilon>0$ we have
$\left|\sum_{j=1}^{m}x_{j}\right|^{1+\varepsilon}\leq m^{\varepsilon}\sum_{j=1}^{m}|x_{j}|^{1+\varepsilon}$. Thus,
@@
Similarly $\mathbb{E}[|U_{k}|^{2(1+\delta)}\mid C_{k}]\leq C(np_{k})^{-(1+\delta)}$. Using the inequality for $|u|\leq 1/2$,
$|\log(1+u)|\leq C(|u|+u^{2})$, we can bound $\mathbb{E}[|\log{\frac{n_{k}}{n}}|^{1+\delta}]$ as below:
@@
As $\ell^{2}=o(n)$, the first term goes to 0. Also the second term goes to 0. Thus,
$\sup_{n}\mathbb{E}[|\log(n_{k}/np_{k})|^{1+\delta}]<\infty$.
@@
We prove this for the $x$ dimension; the proof for the $y$ dimension is identical.
Let $P_{x}^{k}=F_{X|C_{k}}(x_{(i+m_{k})}^{k})-F_{X|C_{k}}(x_{(i-m_{k})}^{k})$. This random variable follows a Beta distribution, $P_{x}^{k}\sim\mathrm{Beta}(2m_{k},n_{k}-2m_{k}+1)$. The conditional mean and variance of its logarithm are given in Johnson:
@@
Let $g(u):=|\log f(u)|^{1+\delta}$ with $\int gf<\infty$. Define the upper step function
$s_{n}(u):=\sum_{k}\big(\sup_{v\in C_{k}}g(v)\big)\mathbf{1}_{C_{k}}(u)$. Since the mesh $\to 0$ and $f$ is continuous with $f>0$ a.e., $s_{n}\downarrow g$ pointwise and $\int s_{n}f\downarrow\int gf<\infty$. Because $(X^{\prime},y_{k}^{*})\in C_{k}$ and
@@
Appendix H Proof Sketch of Proposition 9
The argument is a direct extension of the proof of Proposition 3. We now partition $\mathbb{R}^{d}$ into $\ell^{d}$ hyperrectangles and define the local estimator on each cell $P_{k}$ by
$\hat{f}_{n,\ell}(x_{1},\dots,x_{d})=\frac{n_{k}}{n}\prod_{j=1}^{d}\frac{2m_{k}}{n_{k}\Delta x_{j,i_{j}}^{k}},$
with marginal $m_{k}$-spacings $\Delta x_{j,i_{j}}^{k}$ defined as before. As in Proposition 3, the total mass decomposes into interior and boundary contributions:
$\int_{\mathbb{R}^{d}}\hat{f}_{n,\ell}=\sum_{k=1}^{\ell^{d}}\text{InteriorMass}_{k}+\sum_{k=1}^{\ell^{d}}\text{EdgeMass}_{k}.$
For interior cells, each ratio of grid width satisfies $(\xi_{j,i_{j}+1}^{k}-\xi_{j,i_{j}}^{k})/\Delta x_{j,i_{j}}^{k}=1/(2m_{k})$ by the same telescoping-sum argument as in the bivariate case, yielding
$\text{InteriorMass}_{k}=\frac{n_{k}}{n}(1-1/n_{k})^{d}.$ The boundary terms can again be written in terms of edge factors $E_{j}^{k}$ for $j=1,\dots,d$, each bounded between 0 and $1$. Expanding the product and using $0\leq E_{j}^{k}\leq 1$ shows that
$\text{EdgeMass}_{k}=\frac{n_{k}}{n}\,O\!\left(\frac{m_{k}}{n_{k}}\right),$
with constants depending only on $d$. Summing over $k$ gives
$\int_{\mathbb{R}^{d}}\hat{f}_{n,\ell}=1-O\!\left(\frac{\ell^{d}}{n}\right)+O\!\left(\max_{k}\frac{m_{k}}{n_{k}}\right),$
and the stated convergence follows from $\ell(n)^{d}=o(n)$ and $m_{k}/n_{k}\to 0$.
Appendix I Proof Sketch of Theorem 10
@@
Appendix J Proof Sketch of Theorem 11
@@
$\log\hat{f}_{n,\ell}(X)=\underbrace{\log\tfrac{n_{k}}{np_{k}}}_{\text{(A)}}+\underbrace{\sum_{j=1}^{d}\log S_{j}}_{\text{(B)}}+\underbrace{\sum_{j=1}^{d}\log f(\tilde{X}_{j})}_{\text{(C)}}+\underbrace{(d-1)\log\tfrac{|C_{k}|}{p_{k}}}_{\text{(D)}}$
@@
Code Availability
All code, scripts, and experimental resources necessary to reproduce the results in this paper are publicly available at: https://github.com/hoo0321-design/pss-estimator